# Determinism vs. stochasticity in competitive flour beetle communities

### v1

@@ Abstract
    Additionally, we resolve longstanding issues in flour beetle research by identifying selective egg predation as the mechanism for priority effects, demonstrating the primacy of demographic stochasticity (compared to environmental stochasticity), and reinterpreting classic competition experiments to show that apparent coexistence often represents long-term transient dynamics.
@@ 4 Discussion
Our case study of flour beetles suggests that stochasticity isn’t crucial for understanding coexistence, an idea which finds some support in recent research. One way in which stochasticity can affect coexistence is the storage effect. This fluctuation-dependent coexistence mechanism, wherein species specialize in different states of a variable environment, is thought to be the explanation for how many species can coexist on a few resources (i.e., the paradox of the plankton; Hutchinson,, 1961). However, Stump and Vasseur, (2023) estimate that the storage effect is weaker than classic coexistence mechanisms (e.g., resource partitioning), even in communities where the storage effect is believed to occur; specifically tropical trees of Barro Colorado Island, and communities of annual plants in Spain, California, and Mexico (Godoy and Levine,, 2014; Hallett et al.,, 2017; Zepeda and Martorell,, 2019). The storage effect arises naturally when fecundity fluctuates in a temporally autocorrelated fashion (Schreiber,, 2021), but most animals have temporally uncorrelated or weakly autocorrelated demographic rates (Knape and de Valpine,, 2011; Ferguson et al.,, 2016).
@@
Environmental variation is almost entirely eliminated in our laboratory microcosms, but this may not be as unrealistic as it initially appears. A substantial body of evidence suggests that demographic stochasticity is the primary source of randomness in natural ecosystems (Reed and Hobbs,, 2004; Engen et al.,, 2001; Sæther et al.,, 1998, 2000, 2002; Snyder and Ellner,, 2022), at least on the spatial scales at which most communities are studied. While environmental stochasticity is often favored in analyses due to its mathematical tractability and its faster increase with population density ($n$ vs. $\sqrt{n}$), demographic stochasticity should not be neglected, especially for moderately sized populations typically studied by ecologists. Conceptually, demographic stochasticity is large because it absorbs all individual-level variation that cannot otherwise be explained. Even genetic clones in laboratory conditions can exhibit large differences in life outcomes — crayfish have variable lifespans (Vogt et al.,, 2008) and fish have variable swimming behavior (Bierbach et al.,, 2017; Laskowski et al.,, 2022), putatively caused by a hopelessly intricate interaction between gene expression, microenvironment, and development.

### v2

@@ Abstract
    Additionally, we resolve longstanding issues in flour beetle research by identifying selective egg predation as the mechanism for priority effects, demonstrating the primacy of demographic stochasticity (compared to genetic drift), and reinterpreting classic competition experiments to show that apparent coexistence often represents long-term transient dynamics.
@@ 4 Discussion
Our case study of flour beetles suggests that stochasticity isn’t crucial for understanding coexistence, an idea which finds some support in recent research. One way in which stochasticity can affect coexistence is the storage effect. This fluctuation-dependent coexistence mechanism, wherein species specialize in different states of a variable environment, is thought to be the explanation for how many species can coexist on a few resources (i.e., the paradox of the plankton; Hutchinson,, 1961). However, Stump and Vasseur, (2023) estimate that the storage effect is weaker than classic coexistence mechanisms (e.g., resource partitioning), even in communities where the storage effect is believed to occur, specifically tropical trees of Barro Colorado Island, and communities of annual plants in Spain, California, and Mexico (Godoy and Levine,, 2014; Hallett et al.,, 2017; Zepeda and Martorell,, 2019). The storage effect arises naturally when fecundity fluctuates in a temporally autocorrelated fashion (Schreiber,, 2021), but most animals have temporally uncorrelated or weakly autocorrelated demographic rates (Knape and de Valpine,, 2011; Ferguson et al.,, 2016).
@@
Environmental variation is almost entirely eliminated in our laboratory microcosms, which may raise concerns about generalizability — environmental stochasticity is generally thought to have a larger impact because it increases faster with population density ($n$ vs. $\sqrt{n}$ for demographic stochasticity). However, research suggests that demographic and environmental stochasticity are comparable in significance within an order of magnitude at the spatial scales typical of most studies(Reed and Hobbs,, 2004; Engen et al.,, 2001; Sæther et al.,, 1998, 2000, 2002; Snyder and Ellner,, 2022). Thus, even with reduced environmental variation, the total impact of stochasticity in our experiments remains substantial. Conceptually, demographic stochasticity is large because it absorbs all individual-level variation that cannot otherwise be explained. Even genetic clones in laboratory conditions can exhibit large differences in life outcomes(Vogt et al.,, 2008; Bierbach et al.,, 2017; Laskowski et al.,, 2022), putatively caused by a hopelessly intricate interaction between gene expression, microenvironment, and development.